2/5[15(-10/11)-20=-8(-10/11)+3(-10/11)-6
or
2/5(15p-20)=-8p+3(p-6) and p=-10/11
please help me to solve this. Thank you.

Answer by edjones(8007)

In the first question you are missing a "]" bracket.
.
2/5(15p-20)=-8p+3(p-6)
2(15p-20)=-40p+15(p-6) multiply each side by 5.
30p-40=-40p+15p-90
30p-40=-25p-90
55p=-50
p=-10/11
.
Ed